### v3

# deCIFer: Crystal Structure Prediction from Powder Diffraction Data using Autoregressive Language Models

Abstract
Novel materials drive progress across applications from energy storage to electronics. Automated characterization of material structures with machine learning methods offers a promising strategy for accelerating this key step in material design. In this work, we introduce an autoregressive language model that performs crystal structure prediction (CSP) from powder diffraction data. The presented model, deCIFer, generates crystal structures in the widely used Crystallographic Information File (CIF) format and can be conditioned on powder X-ray diffraction (PXRD) data. Unlike earlier works that primarily rely on high-level descriptors like composition, deCIFer is also able to use diffraction data to perform CSP. We train deCIFer on nearly 2.3M crystal structures and validate on diverse sets of PXRD patterns for characterizing challenging inorganic crystal systems. Qualitative checks and quantitative assessments using the residual weighted profile show that deCIFer produces structures that more accurately match the target diffraction data. Notably, deCIFer can achieve a 94% match rate on test data. deCIFer bridges experimental diffraction data with computational CSP, lending itself as a powerful tool for crystal structure characterization.  [^1] Source code available at: https://github.com/FrederikLizakJohansen/deCIFer
Section 1 Introduction
Characterizing the atomic structure in functional materials is fundamental for understanding and optimizing materials for e.g. new energy technologies. Such characterization can be done using X-ray diffraction (XRD), and in particular, powder X-ray diffraction (PXRD) is widely used in materials chemistry and related fields as a main characterization tool (Cheetham and Goodwin, 2014). Several advances in machine learning (ML) have shown promise in aiding analysis of diffraction data over the last decade (Tatlier, 2011; Bunn et al., 2016; Oviedo et al., 2019; Wang et al., 2020), including recent advances in generative models for crystal structure prediction (CSP) (Jiao et al., 2023; Mohanty et al., 2024; Antunes et al., 2024). These conventional ML-aided CSP methods typically explore the structural phase space guided by high-level descriptors like material composition. Emerging experimental data informed CSP approaches have started to integrate diffraction data directly into the generative process (Kjær et al., 2023; Guo et al., 2024; Riesel et al., 2024; Lai et al., 2025). This shift represents a fundamental departure from conventional CSP.
(a) Overview of the deCIFer model.
(b) Example generations using deCIFer.
Figure 1: (a) Overview of the deCIFer model, which performs autoregressive crystal structure prediction (CSP) from PXRD data, optionally guided by tokenized crystal descriptors. PXRD embeddings are prepended to the CIF token sequence, enabling the generation of structurally consistent CIFs directly from diffraction data. (b) Three examples from the NOMA test set showing deCIFer generations, each illustrating a reference structure, the generated structure and their corresponding PXRD profiles.
In this work, we present deCIFer (Figure 1), a transformer-based model that performs CSP using the rich structural information explicitly represented in experimental data, such as PXRD patterns. Our autoregressive language model directly generates the Crystallographic Information Files (CIFs) – text files that encode for crystal structures. This approach can significantly advance the workflow used for atomic structure characterization based on PXRD methods. Furthermore, we mimic the noise in the PXRD as a result of experimental variations using various transformations to the PXRD patterns.
The performance of deCIFer is evaluated on a diverse set of PXRD patterns, demonstrating its robustness to experimental noise and its ability to generate high-quality crystal structures. deCIFer is the first experimental data informed CSP model capable of generating syntactically correct and structurally meaningful CIFs that accurately reproduce reference PXRD patterns.
Our key contributions in this work are:
  1. 1.
    Integration of experimental PXRD conditioning into an autoregressive transformer for CSP in CIF format; a capability not demonstrated in existing transformer-based generative models.
  2. 2.
    Effective conditioning mechanism for autoregressive models with variable length input.
  3. 3.
    Simulation of realistic experimental settings for PXRD with Gaussian noise and peak broadening.
  4. 4.
    Comprehensive evaluation on two large-scale datasets: NOMA [^2] NOMA stands for NOMAD (Draxl and Scheffler, 2019) OQMD (Kirklin et al., 2015) & M(Jain et al., 2013b) Aggregation. & CHILI-100K (Friis-Jensen et al., 2024), and comparison to state-of-the-art CSP generative models.
Section 2 Background and Related Work
PXRD: It is arguably the most accessible structural probe in solid-state chemistry. Modern benchtop diffractometers can collect a high-quality scan within minutes and are available in most research and industrial laboratories. A PXRD pattern contains diffraction peaks, whose position and intensity contain information on the periodic structure in crystalline materials, i.e. their atomic positions and structure symmetry. Crucially, the forward simulation of PXRD patterns from CIFs is grounded in well-established scattering theory (West, 2014), enabling realistic modelling.
Most quantitative analysis of PXRD data is done using structure refinement, where the parameters in a structural model is refined against experimental data (Young, 1995). This requires a structural starting model of the correct structure type and symmetry. Identifying a model for such analysis is often referred to as fingerprinting, and can be a very challenging task, where chemical intuition as well as extensive database searches are required. Even then, model identification is not always successful, and this hinders further material development.
CSP with LLMs: Transformer architecture (Vaswani et al., 2017) based large language models (LLMs) have seen emerging use in the automation of chemical syntheses (Hocky and White, 2022; Szymanski et al., 2023; M. Bran et al., 2024), data extraction (Gupta et al., 2022; Dagdelen et al., 2024; Polak and Morgan, 2024; Schilling-Wilhelmi et al., 2025), and materials simulation and property prediction (Zhang et al., 2024; Rubungo et al., 2024; Jablonka et al., 2024). However, they are not yet as widely used in tasks like materials design. ChemCrow (M. Bran et al., 2024), for instance, is an LLM-powered chemistry search engine designed to automate reasoning tasks in materials design and other domains.
Recent works have used fine-tuning to adapt LLMs for CSP. Gruver et al. (2024) fine-tuned Llama-2 models (Touvron et al., 2023) on text-encoded atomistic data, enabling tasks like unconditional generation of stable materials. Similarly, Mohanty et al. (2024) fine-tuned LLaMA-3.1-8B (Dubey et al., 2024) using QLoRA (Dettmers et al., 2024) for efficient CIF generation conditioned on material composition and space group. In contrast, the work on CrystaLLM (Antunes et al., 2024) relies on pre-training alone to generate CIFs. It is trained on an extensive corpus of CIFs representing millions of inorganic compounds.
CrystaLLM relies on composition- and symmetry- level crystal descriptors. Its utility, as most other CSP methods, is thus to generate structures without direct reference to experimental data. deCIFer, on the other hand, conditions on PXRD signals, produces structures agreeing with observed diffraction patterns and address the need for models that bridge predictive power with experimental reality.
CSP with diffusion models: In parallel to LLM-based approaches, generative models using diffusion- or flow-based frameworks have also emerged for CSP (Jiao et al., 2023; Miller et al., 2024; Zeni et al., 2025; Xie et al., 2022). These methods rely on composition or partial structural constraints to guide generation, which requires domain knowledge and can limit direct incorporation of experimental data. While they show promise in producing stable crystal configurations, the gap between purely computational crystal structures and experimental data remains a challenge. The recent diffusion-based framework MatterGen (Zeni et al., 2025) addresses some of these challenges by enabling conditioning on a broad range of property constraints, and it can generate structures that are more likely to be synthesizable than previous methods; nevertheless, conditioning on PXRD has not been demonstrated in its current implementation, leaving room for further exploration in integrating experimental data more directly.
Section 3 Methods
Consider a crystal structure in CIF (text input) format that is tokenized into a sequence of length $T_{i}$: ${\mathbf{x}}^{i}=(x^{i}_{1},x^{i}_{2},\dots,x^{i}_{T_{i}})$ (see Appendix A.3 for details on tokenization). The corresponding scattering pattern, which in our case is the PXRD pattern denoted by ${\mathbf{y}}^{i}$, is a continuous-valued vector representing the intensity profile of the scattering pattern. The dataset then consists of pairs of CIFs and PXRD patterns, ${\mathcal{D}}=[({\mathbf{x}}^{i},{\mathbf{y}}^{i})]_{i=1}^{N}$. Given this dataset, we want to minimize the negative conditional log-likelihood over the training data:
${\mathcal{L}}({\mathbf{X}}|{\mathbf{Y}};{\mathbf{\Theta}})=\frac{1}{N}\sum_{i=1}^{N}\left(-\sum_{t=1}^{T_{i}}\log P_{\mathbf{\Theta}}(x^{i}_{t}|x^{i}_{<t},{\mathbf{y}}^{i})\right).$ (1)
This is achieved using a conditional autoregressive model $f_{\mathbf{\Theta}}(\cdot)$ with trainable parameters ${\mathbf{\Theta}}$ based on the transformer architecture (Vaswani et al., 2017). Our autoregressive model, deCIFer, generates structures in CIF format when conditioned with an initial prompt and the PXRD data.
PXRD Conditioning:
The PXRD data can be seen as the structural fingerprint of the structures in CIFs. We use this PXRD data to steer the CSP by using it as a conditioning input to our model.
Following the standard procedure in materials chemistry, for each CIF, we generate the discrete diffraction peak data, given as the set ${\mathcal{P}}=\{(q_{k},i_{k})\}_{k=1}^{n}$ using pymatgen (Ong et al., 2013). During the model training, ${\mathcal{P}}$ is transformed into the PXRD data, ${\mathbf{y}}$, using different simulated experimental conditions. Formally, let ${\mathcal{T}}$ be a set of transformations that can be applied to each ${\mathcal{P}}$.
The family of transformations used in this work are designed to closely mimic the experimental conditions. We define a distribution of transformations ${\mathcal{T}}$ such that each $\tau\sim{\mathcal{T}}$ comprises 1) a random peak broadening with full width at half maximum (FWHM) $\sim\mathcal{U}(0.001,0.100)$ and 2) additive noise with variance $\sigma_{\mathrm{noise}}^{2}\sim\mathcal{U}(0.001,0.050)$. Concretely, each $\tau$ is the composition of these two steps, and new values for broadening and noise are sampled on each draw, ${\mathbf{y}}=\tau({\mathcal{P}})$. Hence for each CIF, the PXRD pattern is transformed slightly differently every time it appears during training. Later for evaluation, we use $\tau_{\mathrm{fixed}}$ with manually specified parameters to test the robustness of the models. The clean transformation $\tau_{0}$ fixes FWHM = $0.05$ and $\sigma_{\mathrm{noise}}^{2}=0$. Examples from ${\mathcal{T}}$ on a PXRD are shown in Figure 9 (in Appendix).
Conditioning Model:
A multi-layered perceptron (MLP) $f_{\mathbf{\Phi}}({\mathbf{y}})$ with trainable parameters ${\mathbf{\Phi}}$ is used to embed the PXRD data into a learnt vector ${\mathbf{e}}=f_{\mathbf{\Phi}}({\mathbf{y}})\in\mathbb{R}^{D}$. This embedding is inserted at the start of the tokenized CIF sequence as a conditioning signal. Joint training of the conditioning network, $f_{\mathbf{\Phi}}$, and the autoregressive model, $f_{\mathbf{\Theta}}$, enables the integration of structural information from the PXRD profiles during the generation of CIFs. This joint training along with transformations to the PXRD data results in the final training objective ${\mathcal{L}}({\mathbf{X}}|{\mathbf{Y}};{\mathbf{\Theta}},{\mathbf{\Phi}})$.
Sequence Packing and Isolating CIFs:
To effectively train on variable-length CIF sequences, we implement a batching strategy during training, inspired by recent sequence-packing methods that prevent cross-contamination and improve throughput (Kosec et al., 2021). Our approach concatenates multiple tokenized CIF sequences, each of length $T_{i}$ (excluding the conditioning embedding) into segments of fixed context length, $C$. Formally, consider a sequence $\mathbf{S}=[\mathbf{e}^{1},{\mathbf{t}}^{1}_{1},\dots,{\mathbf{t}}^{1}_{T_{1}},\mathbf{e}^{2},{\mathbf{t}}^{2}_{1},\dots,{\mathbf{t}}^{n}_{k}]$, where $n$ is the last fully or partially included CIF. In this notation ${\mathbf{e}}^{i}$ denotes a $D$-dimensional conditioning embedding for the $i$-th CIF, while ${\mathbf{t}}_{j}^{i}$ is the $D$-dimensional input embedding for $x_{j}^{i}$, the $j$-th token of the $i$-th CIF. We choose $k$ such that $|\mathbf{S}|=C$. If adding another CIF exceeds $C$, that CIF is split at the boundary and continued in the next segment. In practice, we set $C=3076$ based on the available GPU memory to balance throughput and memory constraints, which exceeds the length of the longest tokenized CIFs in the NOMA dataset. Although our method guarantees efficient batch utilization, it does occasionally split exceptionally long CIFs between batches ($\approx\!0.04$% of samples in the NOMA dataset; see Figure 10 in the Appendix). To mitigate this, we shuffle the training set at the start of each epoch so that previously split CIFs are more likely to appear in full in subsequent mini-batches, allowing the model to learn from complete sequences.
To isolate different CIFs in the same sequence, we employ an attention mask ${\mathbf{M}}$, defined such that $M_{kl}=1$ iff tokens $k$ and $l$ belong to the same CIF, and $M_{kl}=0$ otherwise. This results in a block-wise diagonal, upper-triangular, attention matrix as shown in Figure 8 (Appendix). To prevent positional information from leaking across CIF boundaries we also reset positional encodings at the start of each CIF by assigning positions from 00 to $T_{i}-1$ for the $i$-th sequence.
Section 4 Dataset and Experiments
Dataset:
We use two large-scale public datasets here. The first, NOMA, is a large-scale compilation of crystal structures curated in CrystaLLM (Antunes et al., 2024), which draws from the Materials Project (April 2022) (Jain et al., 2013b), OQMD (v. 1.5, October 2023) (Kirklin et al., 2015), and NOMAD (April 2023) (Draxl and Scheffler, 2019) databases. The second, CHILI-100K (Friis-Jensen et al., 2024), is a large-scale dataset of experimentally determined crystal structures obtained from a curated subset of the Crystallography Open Database (COD) (Gražulis et al., 2009). We employ NOMA for both training and testing our models, whereas CHILI-100K is used exclusively for testing. Both datasets are open-source and available for download. [^3] NOMA: github.com/lantunes/CrystaLLM (CC-BY 4.0 licence), CHILI-100K: github.com/UlrikFriisJensen/CHILI (Apache 2.0 licence).
Preprocessing:
We perform standard data preprocessing steps outlined in CrystaLLM before performing additional curation to address complexities such as ionic charges and to ensure consistent structural representation between NOMA and CHILI-100K. The resulting NOMA dataset comprises approximately $2.3$M CIFs, spanning 111110101010 elements, including elements up to atomic number $94$ (excluding polonium, astatine, radon, francium, and radium). Duplicate structures were filtered by selecting the configuration with the lowest volume per formula unit. All structures were converted to a standardized CIF format using the pymatgen library (Ong et al., 2013) and all floating point values were rounded to four decimal places. The resulting CHILI-100K dataset consists of $\approx\!8.2$K CIFs, spanning 11118888 elements, including elements up to atomic number $85$. Figures 10 and 11 (in Appendix) show the distribution of elemental compositions, space groups, and more attributes for the two datasets.
Additionaly, a known challenge in datasets like NOMA is the disproportionate representation of high-symmetry structures, which can skew model learning and evaluation (Davariashtiyani et al., 2024; Zhang et al., 2023). To mitigate this, we stratify the dataset during the train/validation/test split based on space group labels. For details on stratification, see Section A.9.
Finally, CIFs from both datasets were tokenized into a vocabulary of $373$ tokens, encompassing CIF tags, space group symbols, element symbols, numeric digits, punctuation marks, padding tokens and a special conditioning token. For a detailed explanation of the standardization and tokenization processes, refer to Section A.2 and A.3 in the Appendix.
Model Hyperparameters:
deCIFer has two learnable components, $f_{{\mathbf{\Phi}}}$ and $f_{{\mathbf{\Theta}}}$. The encoder $f_{{\mathbf{\Phi}}}$ is a 2-layer MLP that takes a PXRD profile of dimension $1000$ and outputs a $512$-dimensional embedding. This embedding is prepended as the conditioning-token to the sequence of CIF tokens, where each CIF token is also of $D=512$ dimensions. $f_{{\mathbf{\Theta}}}$ is a decoder-only transformer (Vaswani et al., 2017) with $8$ layers, each containing $8$ attention heads. We use a context length of $3076$ and a batch size of $32$. A linear warm-up of the learning rate is applied over the first $100$ epochs, followed by a cosine decay schedule using AdamW (Loshchilov and Hutter, 2017) ($\beta_{1}=0.9,\beta_{2}=0.95$, weight decay $10^{-1}$) for $50$K epochs with gradient accumulation of $40$ steps on a single NVIDIA A100 GPU with mixed-precision acceleration. All aspects of the model architecture were implemented in Pytorch (Paszke et al., 2019). $f_{{\mathbf{\Phi}}}$ has $\approx\!0.78$M trainable parameters and $f_{{\mathbf{\Theta}}}$ contains $\approx\!26.94$M, resulting in a deCIFer model with $27.72$M parameters. The code is open-source and additional details can be found in Section A.9 in the Appendix.
Evaluation:
Figure 2 summarizes the evaluation pipeline which shows how a reference CIF from the test set is used to generate discrete peaks, $\mathcal{P}=\{(q_{k},i_{k})\}_{k=1}^{n}$, which are transformed into a PXRD profile $\mathbf{y}=\tau_{\mathrm{fixed}}(\mathcal{P})$, where $\tau_{\mathrm{fixed}}$ is selected according to the desired experimental setting. The resulting PXRD, along with any optional crystal descriptors, is tokenized and passed to deCIFer to produce a new CIF ($\text{CIF}^{*}$). For evaluation, the generated structures are compared using three common metrics. Here a clean reference transformation $\tau_{0}$ (FWHM$=0.05$, $\sigma_{\mathrm{noise}^{2}}=0$) is used.
Figure 2: Evaluation pipeline: A test set CIF generates a PXRD profile, tokenized for deCIFer to produce a new CIF, compared to the reference using a clean transformation.
1) Residual weighted profile ($R_{\mathrm{wp}}$) is applied to convolved continuous PXRD profiles, measuring the discrepancy between a reference PXRD profile, ${\mathbf{y}}$, and a generated PXRD profile, ${\mathbf{y}}^{*}$. Formally: $R_{\mathrm{wp}}=\sqrt{\frac{\sum_{i}w_{i}(y_{i}-y_{i}^{*})^{2}}{\sum_{i}w_{i}y_{i}^{2}}}$ with all weights $w_{i}=1$ here, following convention.
2) Match rate (MR) uses StructureMatcher (Ong et al., 2013) to assess structural similarity between reference and generated CIFs. Two structures are considered a match if their lattice parameters, atomic coordinates, and space group symmetries are within the defined tolerances. MR is the fraction of matching structures. See Section A.7 in the Appendix for more information.
3) Validity (Val.) ensures the internal consistency of each generated CIF by checking formula consistency, site multiplicity, bond lengths, and space group alignment. A CIF is deemed overall valid only if it passes all four checks. Detailed explanations of these metrics are provided in Section A.6 in the Appendix.
Experiments:
We conducted a series of experiments to evaluate deCIFer’s ability to perform CSP when conditioned on PXRD profiles, space group, and composition. First, we compare against three recent state-of-the-art CSP models. Next, we compare the baseline performance of deCIFer against an unconditioned variant (U-deCIFer) to assess the direct impact of PXRD conditioning on the generated structures. Next, we introduce controlled noise and peak broadening into the input PXRD data, examining deCIFer’s robustness under more challenging scenarios resembling real-world PXRD data. Finally, we apply deCIFer (trained on NOMA) to CHILI-100K (Friis-Jensen et al., 2024) to demonstrate its scalability and performance on more complex crystal systems. In these experiments, we generate one CIF for each reference sample in the test sets. For consistent evaluation, each reference CIF is also processed through the fixed, clean transformation $\tau_{0}$ (FWHM $=0.05$, $\sigma_{\mathrm{noise}}^{2}=0$).
For some experiments we allow space group and composition to be specified as crystal descriptors. Since CIFs inherently encode space group and composition as text, we treat these descriptors as standard tokens in our vocabulary (see Section A.3 in the Appendix). The space group appears in the CIF header, while composition is stored as element–count pairs. During inference, these tokens are optionally inserted into the CIF.
Section 5 Results
Baseline Comparisons with State-of-the-art:
To evaluate the structure generation quality of deCIFer, we compare against three recent state-of-the-art CSP models: CDVAE (Xie et al., 2022), DiffCSP (Jiao et al., 2023), and CrystaLLM (Antunes et al., 2024). Following the single-sample evaluation setup used in prior work (Antunes et al., 2024), we generate one structure per composition and assess accuracy using two metrics: match rate and root-mean-square error (RMSE) in atomic positions (Å). To ensure a fair and consistent evaluation, we adopt the official train-test splits provided by DiffCSP (Jiao et al., 2023) for all four benchmark datasets: Perov-5 (Castelli et al., 2012a; b), Carbon-24 (Pickard, 2020), MP-20 (Jain et al., 2013a), and MPTS-52 (Baird, 2023).
As shown in Table 1, deCIFer outperforms all baselines on Perov-5 and Carbon-24, highlighting the benefit of PXRD conditioning when the diffraction data are informative. However, on MP-20 and MPTS-52, CrystaLLM-large outperforms deCIFer, likely due to its exclusive emphasis on composition-structure mappings, which align closely with the test set. This highlights a key trade-off in experimental data informed CSP: while conditioning can enhance structure generation in complex cases, it can also introduce ambiguity when the PXRD signal is weak or the model’s training priors are strong.
Table 1: Performance comparison on four public CSP benchmarks: Perov-5 (Castelli et al., 2012a; b), Carbon-24 (Pickard, 2020), MP-20 (Jain et al., 2013a), and MPTS-52 (Baird, 2023). Following the single-sample evaluation protocol used in prior work (Antunes et al., 2024), one structure is generated per test composition. We report the match rate (%) based on structural equivalence under StructureMatcher and the atomic root-mean-square error (RMSE in Å) after alignment.
Perov-5 Carbon-24 MP-20 MPTS-52
[…]
CDVAE 45.31 0.1138 17.09 0.2969 33.90 0.1045 5.34 0.2106
DiffCSP 52.02 0.0760 17.54 0.2759 51.49 0.0631 12.19 0.1786
CrystaLLM-small 47.95 0.0966 21.13 0.1687 55.85 0.0437 17.47 0.1113
CrystaLLM-large 46.10 0.0953 20.25 0.1761 58.70 0.0408 19.21 0.1110
U-deCIFer 50.55 0.1177 17.33 0.1526 44.99 0.0784 11.80 0.1563
deCIFer 85.29 0.0491 37.16 0.1970 43.51 0.0763 11.44 0.1346
Importance of PXRD Conditioning:
We compare deCIFer and U-deCIFer to study the effect of PXRD conditioning. We evaluate each model in three settings: without any crystal descriptors (“none”), with only compositional information (“comp.”), and with both compositional and space group information (“comp. + s.g.”). Both models share the same architectural backbone, but deCIFer receives structural guidance via the PXRD conditioning.
Figure 3 shows that deCIFer achieves lower average $R_{\mathrm{wp}}$ than U-deCIFer across all descriptor settings. Figure 3 visualizes the distributions of $R_{\mathrm{wp}}$ for both models, highlighting the large performance improvement due to PXRD conditioning, and a smaller but still notable improvement due to composition conditioning. Figure 4 shows that structures of more common crystal systems lead to better outcomes, while under-represented, lower symmetry crystal systems remain challenging. The figure further illustrates the effectiveness of PXRD conditioning with three examples from the NOMA test set, showcasing the range of generated structures, from near-perfect alignment to structural mismatch. While U-deCIFer benefits from the addition of crystal descriptors, it never reaches the accuracy of deCIFer. Overall, these results demonstrate that conditioning on PXRD data and crystal descriptors substantially enhances the model’s ability to generate CIFs with close alignment to desired PXRD profiles.
Desc. Model $R_{wp}$ $(\mu\pm\sigma)\downarrow$ Val. (%) $\uparrow$ MR (%) $\uparrow$
none U-deCIFer $1.24$$\pm 0.26$ 93.49 0.00
deCIFer $0.32$$\pm 0.34$ 92.66 5.01
comp. U-deCIFer $0.82$$\pm 0.41$ 93.78 49.30
deCIFer $0.25$$\pm 0.29$ 93.73 91.50
comp.+s.g. U-deCIFer $0.65$$\pm 0.36$ 93.72 87.07
deCIFer $0.24$$\pm 0.29$ 93.90 94.53
Figure 3: Left: Distribution of $R_{\mathrm{wp}}$ for deCIFer and U-deCIFer on the NOMA test set with boxplots. Lower $R_{\mathrm{wp}}$ indicates better CIF alignment. Right: Performance for 20K NOMA test samples using deCIFer and U-deCIFer with different descriptors: none (no descriptors), comp. (composition), and comp.+ s.g. (composition + space group). Metrics include validity (Val.) and match rate (MR).
Figure 4: Left: Average $R_{\mathrm{wp}}$ by crystal system for deCIFer on the NOMA test set shows better performance for common high symmetry systems and higher $R_{\mathrm{wp}}$ for rare low symmetry systems. Right: Examples from the NOMA test set highlight this trend with predicted structures from PXRD and composition maintaining reasonable matches even for low symmetry systems with higher $R_{\mathrm{wp}}$.
Robustness to Perturbations in PXRD Conditioning
We evaluated deCIFer’s ability to generate accurate CIFs under varying levels of additive noise and peak broadening in the PXRD conditioning input, while also providing the composition as a crystal descriptor. Building upon the baseline scenario of clean, noise-free PXRD data, we tested several increasingly challenging conditions: maximum noise, maximum peak broadening, combined noise and broadening, and out-of-distribution (OOD) broadening levels beyond the model’s training range.
Figure 5 (and Figure 16 in Appendix) show that while additive noise results in slight performance degradation, the model remains robust to in-distribution noise and peak broadening. We also observe moderate performance degradation with OOD broadening or significantly higher noise levels. Unsurprisingly, we observe that lower-symmetry crystal systems remain more challenging to predict under perturbed conditions across all crystal systems. This is depicted in Figure 14 in the Appendix.
OOD Evaluation on CHILI-100K
To evaluate the generalization of deCIFer to more complex crystal systems, we tested its performance on the CHILI-100K dataset (Friis-Jensen et al., 2024) which had no overlap with the NOMA training data for deCIFer. Unlike synthetic datasets, CHILI-100K presents a closer approximation to real-world challenges, including complex structural motifs and a broader distribution of crystal symmetries. Additionally, CHILI-100K contains a significantly higher proportion of lower-symmetry structures compared to synthetic datasets like NOMA (see the sample distribution in Appendix Figure 15).
The results on CHILI-100k dataset are summarized in Figure 5. deCIFer maintains a reasonable level of structural accuracy on this challenging dataset. The relatively low validity score is partly due to challenges with bond length validity, which was notably lower than the other validity metrics. A full breakdown of the validity metrics for this experiment is available in Section A.6 in the Appendix.
Despite the performance drop, deCIFer demonstrated robustness to added noise and peak broadening in PXRD inputs, with stable $R_{\mathrm{wp}}$ values across perturbed conditions as seen in Figure 5. This stability, along with its ability to generate a variety of structural features, suggests that deCIFer could be useful for practical applications involving experimentally derived PXRD data.
Table 2: Performance on CHILI-100K without (U-deCIFer) and with (deCIFer) PXRD conditioning.
Model $R_{\mathrm{wp}}\downarrow$ Val. (%$\uparrow$ MR (%$\uparrow$
U-deCIFer $0.96\pm 0.32$ $43.26$ $25.92$
deCIFer $0.70\pm 0.37$ $41.83$ $37.34$
Dataset ($\sigma_{\mathrm{noise}}^{2}$, FWHM) $R_{\mathrm{wp}}\;(\mu\pm\sigma)\downarrow$ Val. (%)\;$\uparrow$ MR (%)\;$\uparrow$
NOMA ID: (0.00, 0.05) 0.25$\pm$0.29 93.73 91.50
ID: (0.05, 0.10) 0.31$\pm$0.30 93.77 89.28
OOD: (0.10, 0.20) 0.65$\pm$0.34 91.66 77.66
CHILI-100K ID: (0.00, 0.05) 0.70$\pm$0.37 41.83 37.34
ID: (0.05, 0.10) 0.73$\pm$0.36 40.95 35.97
OOD: (0.10, 0.20) 0.87$\pm$0.33 33.62 26.09
Figure 5: Left: Distribution of $R_{\mathrm{wp}}$ for deCIFer on NOMA and CHILI-100K using PXRD conditioning and composition (comp.) across three scenarios: clean, high noise/broadening, and out-of-distribution (OOD), with noise and FWHM values indicated. Right: Corresponding table of performance metrics, including $R_{\mathrm{wp}}$, overall validity (Val.), and match rate (MR) for each scenario.
To assess the impact of PXRD conditioning on OOD data, we compare deCIFer with U-deCIFer on the CHILI-100K test set. Results are shown in Table 2. We observe that PXRD conditioning lowers $R_{\mathrm{wp}}$ and has a significant positive effect on the match rate, showcasing that the conditioning drives improvement and not merely compositional priors.
Section 6 Discussion and Outlook
PXRD-driven Structure Generation: The experiments on NOMA and CHILI-100K in Section 5 clearly show the utility of including conditioning data such as PXRD when performing CSP. deCIFer consistently achieves strong performance on the NOMA test set, where the diversity of crystal systems and the availability of high-quality PXRD signals allow the model to accurately generate structures that align with the diffraction data. This highlights a key advantage of PXRD conditioning: when the diffraction data are clear and informative, deCIFer can use them to resolve complex atomic arrangements, even in cases where composition-only models struggle.
However, the results also make it clear that deCIFer, like any PXRD-conditioned model, cannot be used as a black-box end-to-end solver. The model’s ability to generate accurate structures is tied to the quality and relevance of the PXRD signal. On datasets like MP-20, where learned composition–structure priors dominate, PXRD conditioning can sometimes interfere with accurate generation by introducing conflicting guidance. This tension between learned priors and PXRD conditioning is not unique to deCIFer but reflects a broader challenge in PXRD-informed CSP.
By embedding the PXRD data, ${\mathbf{y}}$, into a learnable conditional embedding, ${\mathbf{e}}=f_{\mathbf{\Phi}}({\mathbf{y}})$, and prepending it onto prompts for deCIFer, we have demonstrated a general paradigm for integrating additional conditional data in materials design, an idea which can be extended to the incorporation of other material properties. In cases where multiple experimental data sources are available, they can be easily injected into the generative process using specific conditioning models. That is, if $P$ properties are available such that ${\mathbf{y}}=\{{\mathbf{y}}_{1},\dots,{\mathbf{y}}_{P}\}$, then deCIFer can be extended to incorporate these additional data using additional conditioning models, $f_{{\mathbf{\Phi}}_{1}},\dots,f_{{\mathbf{\Phi}}_{P}}$, that can be jointly trained with the generative model, $f_{\mathbf{\Theta}}$, via a training objective of the form ${\mathcal{L}}({\mathbf{X}}|{\mathbf{Y}}_{1},\dots,{\mathbf{Y}}_{P};{\mathbf{\Theta}},{\mathbf{\Phi}}_{1},\dots,{\mathbf{\Phi}}_{P})$.
Consistency in CIF generation: To further investigate the consistency and variability of deCIFer, we generated 16K CIFs for the same PXRD profile with different crystal descriptors (“none”, “comp.”, and “comp+s.g.”). Figure 6 illustrates our findings for a challenging monoclinic crystal system of Sr2Cd2Se4 from the NOMA test set. When unconstrained by crystal descriptors, the model generates a wide diversity of cell-parameters, compositions, and space groups, yet the $R_{\mathrm{wp}}$ values tend to cluster in a relatively narrow range. In contrast, imposing compositional and, especially, space group constraints yields much tighter cell-parameter distributions and a broader $R_{\mathrm{wp}}$ distribution. The broader distributions highlight how the $R_{\mathrm{wp}}$ metric is highly sensitive to even small structural deviations and shows the importance of complementing with other validation methods. Despite these variations, the overall structural match rate to the reference CIF remains high. If the composition or space group is known with high confidence, incorporating these constraints can speed up convergence to a more accurate structural solution. Conversely, if the objective is to explore the material landscape more broadly, unconstrained generation with PXRD can be advantageous.
Figure 6: deCIFer-sampled structures for a monoclinic Sr2Cd2Se4444 PXRD profile (16K samples). a) Reference structure. b) Distribution of $R{\mathrm{wp}}$ for generated CIFs. c) Examples of generated structures showing best, median, and diverse samples. d) Distribution of sampled crystal systems. e) Histograms of cell lengths (a, b, c) and angles ($\alpha,\beta,\gamma$) with reference values as dotted lines.
Limitations: In materials science, the concept of data leakage is nuanced due to the challenges in defining novelty and diversity, as emphasized in recent discussions on scaling AI for materials discovery (Cheetham and Seshadri, 2024). Although we stratify the NOMA dataset by space group bins and use CHILI-100K exclusively for testing, there remains the possibility of implicit overlaps, such as structurally similar crystal entries or shared compositional biases.
Furthermore, the vast size and diversity of the NOMA dataset as well as the thorough preprocessing steps (e.g., de-duplication, filtering, and standardization), coupled with the independent curation of CHILI-100K, significantly mitigate the impact of any potential overlaps.
While PXRD conditioning enhances structure prediction by introducing rich physical constraints, it also inherits a well-known limitation: the inability to distinguish between homometric structures. These are distinct atomic arrangements that yield indistinguishable diffraction patterns (Patterson, 1944). This ambiguity has been documented in real structural analyses (Schneider et al., 2010). Because metrics such as $R_{\mathrm{wp}}$ operate only on the diffraction pattern, they cannot capture homometric degeneracy. Nevertheless, several strategies can help mitigate this limitation (Shen et al., 2022). Particularly, our experiments with deCIFer show that even partial inclusion of complementary information, such as composition, can mitigate near-degeneracy cases where the only differences in the PXRD are subtle variations in relative peak intensities.
Section 7 Conclusion
In this work, we introduced deCIFer as a data-informed approach to CSP. deCIFer is an autoregressive transformer-based language model that generates CIFs directly from PXRD data developed using lab-scale compute resources with a single GPU. By conditioning on simulated PXRD profiles, deCIFer captures fine-grained structural information beyond what is possible with composition- or symmetry-based information alone. Evaluations on large synthetic and experimentally derived datasets highlight deCIFer’s robust performance in scenarios with varying levels of noise and peak broadening.
Ultimately, deCIFer represents a step forward in PXRD-informed CSP, bridging the gap between purely composition-based generative models and structure generation directly guided by experimental data. It is most useful when the diffraction signal is informative and when combined with careful downstream verification. For real-world use, this balance must be clearly understood, and model predictions must be treated as high-quality starting points rather than definitive solutions.
Acknowledgments
The authors thank Richard Michael and Adam F. Sapnik for useful feedback.
This work is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No. 804066). We are grateful for funding from University of Copenhagen through the Data+ program. RM acknowledges funding provided by the Wallenberg AI, Autonomous Systems, and Software Program (WASP), supported by the Knut and Alice Wallenberg Foundation.
Appendix A Appendix
Section A.1 Code and Data Availability
The code for training and using the deCIFer model is open source and released under the MIT License. Official source code is available here: https://github.com/FrederikLizakJohansen/deCIFer.
Section A.2 CIF Syntax Standardization
To enhance the transformer model to process CIFs effectively, we standardized all CIFs in the dataset. Inspired by CrystaLLM [Antunes et al., 2024], we employed similar pre-processing and tokenization strategies, incorporating additional steps to ensure that CHILI-100K [Friis-Jensen et al., 2024] was aligned to the standardized format of NOMA, by the removal certain details such as oxidation states and partial occupancies. We employ the following steps:
  1. 1.
    Uniform Structure Conversion: CIFs were converted to pymatgen.Structure [Ong et al., 2013] objects to provide a consistent base representation.
  2. 2.
    Standardized CIF Regeneration: Using pymatgen.CifWriter [Ong et al., 2013], CIFs were regenerated to ensure uniform formatting, eliminate customs headers, etc.
  3. 3.
    Data Tag Normalization: The reduced formula, following the data_ tag was replaced with the full cell composition, sorted by atomic number for consistency.
  4. 4.
    Symmetry Operator Removal: Symmetry operators were excluded during pre-processing to simplify the data, but reintroduced during evaluation for validating structural matches. This can easily be done because the reintroduction process uses the space group information retained in the pre-processed files, ensuring compatibility and accuracy.
  5. 5.
    Incorporation of Extra Information: Custom properties that are easily derived from the composition of each CIF, such as electronegativity, atomic radius, and covalent radius, were appended to each CIF to maximize the readily available information within each CIF.
  6. 6.
    Oxidation State and Occupancy Filtering: Oxidation state refers to the charge of an atom within a compound, which can vary depending on chemical bonding. Occupancy indicates the fraction of a particular atomic site that is occupied in the crystal structure (e.g., a value of 1.0 represents a fully occupied site, while 0.5 indicates partial occupancy). All traces of oxidation states were removed, and only crystal structures with full occupancy were retained. This ensures consistency by aligning CHILI-100K [Friis-Jensen et al., 2024] with the standardized format of NOMA [Antunes et al., 2024].
  7. 7.
    Numerical Value Normalization: Numerical values were rounded to four decimal places.
Figure 7 shows a pre-processed and standardized CIF from the NOMA dataset alongside its corresponding unit cell representation and a realisation of its corresponding PXRD profile, as could be input into deCIFer.
Figure 7: Illustration of a CIF after applying the pre-processing and standardization steps described. Also shown are the corresponding unit cell representation using VESTA [Momma and Izumi, 2008] for visualization and the simulated PXRD profile (with $\sigma^{2}=0.00$ and FWHM=0.01). The red highlight in the CIF indicates where the original symmetry operators were replaced during pre-processing and would be restored for evaluation.
Section A.3 CIF Tokenization
To process CIF files effectively, we tokenized each file into a sequence of tokens using a custom vocabulary tailored to crystallographic data in the CIF format. Each CIF was parsed to extract key structural and chemical information, such as lattice parameters, atomic positions, and space group symbols. Numerical values were tokenized digit-by-digit, including decimal points and special characters as separate tokens. Table 3 shows all supported tokens.
Table 3: Supported atoms, CIF tags, space groups, numbers, and special tokens.
Category Num. Tokens
Atoms 89 Si C Pb I Br Cl Eu O Fe Sb In S N U Mn Lu Se Tl Hf Ir Ca Ta Cr K Pm Mg Zn Cu Sn Ti B W P H Pd As Co Np Tc Hg Pu Al Tm Tb Ho Nb Ge Zr Cd V Sr Ni Rh Th Na Ru La Re Y Er Ce Pt Ga Li Cs F Ba Te Mo Gd Pr Bi Sc Ag Rb Dy Yb Nd Au Os Pa Sm Be Ac Xe Kr He Ne Ar
CIF Tags 31 data_
loop_
_symmetry_space_group_name_H-M
_symmetry_Int_Tables_number
_cell_length_a
_cell_length_b
_cell_length_c
_cell_angle_alpha
_cell_angle_beta
_cell_angle_gamma
_cell_volume
_atom_site_fract_x
_atom_site_fract_y
_atom_site_fract_z
_atom_site_occupancy
_symmetry_equiv_pos_as_xyz
_chemical_formula_structural
_cell_formula_units_Z
_chemical_name_systematic
_chemical_formula_sum
_atom_site_symmetry_multiplicity
_atom_site_attached_hydrogens
_atom_site_label
_atom_site_type_symbol
_atom_site_B_iso_or_equiv
_symmetry_equiv_pos_site_id
_atom_type_symbol
_atom_type_electronegativity
_atom_type_radius
_atom_type_ionic_radius
_atom_type_oxidation_number
Space Groups 230 P6/mmm Imma P4_32_12 P4_2/mnm Fd-3m P3m1 P-3 P4mm P4_332 P4/nnc P2_12_12 Pnn2 Pbcn P4_2/n Cm R3m Cmce Aea2 P-42_1m P-42m P2_13 R-3 Fm-3 Cmm2 Pn-3n P6/mcc P-6m2 P3_2 P-3m1 P3_212 I23 P-62m P4_2nm Pma2 Pmma I-42m P-31c Pa-3 Pmmn Pmmm P4_2/ncm I4/mcm I-4m2 P3_1 Pcc2 Cmcm I222 Fddd P312 Cccm P6_1 F-43c P6_322 Pm-3 P3_121 P6_4 Ia-3d Pm-3m P2_1/c C222_1 Pc P4/n Pba2 Ama2 Pbcm P31m Pcca P222 P-43n Pccm P6_422 F23 P42_12 C222 Pnnn P6_3cm P4_12_12 P6/m Fmm2 I4_1/a P4/mbm Pmn2_1 P4_2bc P4_22_12 I-43d I4/m P4bm Fdd2 P3 P6_122 Pnc2 P4_2/mcm P4_122 Cmc2_1 P-6c2 R32 P4_1 P4_232 Pnna P422 Pban Cc I4_122 P6_3/m P6_3mc I4_1/amd P4_2 P4/nmm Pmna P4/m Fm-3m P4/mmm Imm2 P4/ncc P-62c Ima2 P6_5 P2/c P4/nbm Ibam P6_522 P6_3/mmc I4/mmm Fmmm P2/m P-4b2 I-4 C2/m P4_2/mmc P4 Fd-3c P4_3 P2_1/m I-43m P-42c F4_132 Pm Pccn P-4n2 P4_132 P23 I4cm R3c Amm2 Immm Iba2 I4 Fd-3 P1 Pbam P4_2/nbc Im-3 P4_2/nnm Pmc2_1 P-31m R-3m Ia-3 P622 F222 P2 P-1 Pmm2 P-4 Aem2 P6_222 P-3c1 P4_322 I422 Pnma P6_3 P3c1 Pn-3 P4nc P-6 P4/mcc I2_12_12_1 P4_2/mbc P31c Ccc2 P4_2/nmc P6_3/mcm C2 Pbca P-4c2 I4_1cd P2_1 P3_112 P4_2mc Pn-3m C2/c R3 P-43m I432 P222_1 I-42d I-4c2 P6cc P6_2 P3_221 P321 Pca2_1 I4_1/acd I4_132 F432 Pna2_1 Ccce Ibca P4/mnc I4_1md P2_12_12_1 R-3c I2_13 P-4m2 Pm-3n I4mm F-43m Pnnm P-42_1c Cmmm P6mm P4_2cm P4_2/m Im-3m Fm-3c I4_1 P4cc Cmme
Numbers 10 1 2 3 4 5 6 7 8 9 0
[…]
Section A.4 Attention Masking Strategy
Figure 8 provides a detailed visualization of the attention masking strategy employed in our model. It illustrates the log-mean attention weights (averaged over all heads) for a sample sequence, highlighting the isolation of CIFs through attention masking. The figure also demonstrates how the embeddings of the second CIF attend to the conditioning PXRD embedding. Lighter shades in the figure correspond to stronger attention values.
Figure 8: Visualization of the attention masking strategy, showing the log-mean attention weights (averaged over all heads) for an example sequence and highlighting how CIFs are isolated using the attention mask. The figure also illustrates how the embeddings of the second CIF attend to the conditioning PXRD embedding. Lighter shades indicate stronger attention.
Section A.5 PXRD Simulation
What do the axes in PXRD mean?
In a typical PXRD experiment, the x-axis corresponds to the magnitude of the scattering vector, commonly denoted by $Q$ (in units of Å-1), or sometimes the diffraction angle $2\theta$. In this work, we use $Q=\tfrac{4\pi\sin\theta}{\lambda}$ where $\lambda$ is the radiation wavelength and $\theta$ is the scattering angle. The y-axis represents the scattered intensity observed at each $Q$-value, sometimes normalized to have a maximum intensity of $1$.
Figure 9: Simulated PXRD profiles with fixed transformation of FWHM and $\sigma_{\mathrm{noise}}^{2}$ as indicated. Discrete peaks, ${\mathcal{P}}=\{(q_{k},i_{k})\}_{k=1}^{n}$, are shown in red, while the convolved PXRD profiles, ${\mathbf{y}}$, are shown in blue. Examples with minimal and maximal noise and broadening levels are shown for a compound with composition CdRhBr2 and space group R3m.
Peak data and transformations.
Following the methods section, we start with the discrete diffraction peak data: ${\mathcal{P}}={(q_{k},i_{k})_{k=1}^{n}}$, where each $q_{k}$ is the center of a reflection peak, and $i_{k}$ is the associated peak intensity. To simulate experimental effects, we apply transformation $\tau\sim{\mathcal{T}}$, which includes peak broadening and additive noise.
Peak broadening.
For each peak $k$, centered at $q_{k}$, we convolve an idealized delta function peak with a pseudo-Voigt profile. At the continuous variable $Q$, the psuedo-Voigt profile is the mixture of a Lorentzian $L$ and a Gaussian $G$, such that
$\mathrm{PV}_{k}(Q-q_{k})=\eta L(Q-q_{k})+(1-\eta)G(Q-q_{k}),$ (2)
where $0\leq\eta\leq 1$ is fixed at $\eta=0.5$ in this work.
Let $\mathrm{FWHM}$ denote the full width at half maximum. The Lorentzian half-width is then $\gamma=\tfrac{\mathrm{FWHM}}{2}$, making
$L(Q-q_{k})=\frac{1}{1+\left(\tfrac{Q-q_{k}}{\gamma}\right)}.$ (3)
The Gaussian standard deviation is $\sigma=\tfrac{\mathrm{FWHM}}{2\sqrt{2\ln 2}}$, making
$G(Q-q_{k})=\exp\left(-\tfrac{1}{2}\left(\tfrac{Q-q_{k}}{\sigma}\right)^{2}\right).$ (4)
Convolved PXRD.
Given the peak centers $q_{k}$, itensities $i_{k}$, and a choice of FWHM, we obtain the convolved PXRD profile
$I_{\mathrm{conv}}(Q)=\sum_{k=1}^{n}i_{k}\;\mathrm{PV}_{k}(Q-q_{k}).$ (5)
Afterwards, we normalize $I_{\mathrm{conv}}(Q)$ so that its maximum intensity is $1$.
Noise addition.
Let $\epsilon(Q)$ be drawn from a zero-mean Gaussian distribution with variance $\sigma^{2}_{\mathrm{noise}}$. This yields the final transformed intensity PXRD profile:
$I(Q)=I_{\mathrm{conv}}(Q)+\epsilon(Q).$ (6)
Implementation details.
In practice, we use the XRDCalculator from the pymatgen library [Ong et al., 2013] for generating the initial discrete peak data ${\mathcal{P}}$. For training, we sample $Q$-values in $[Q_{\mathrm{min}},Q_{\mathrm{max}}]$ at increments of $Q_{\mathrm{step}}$. We then apply random transformations $\tau$ during model training. Specific parameters for FWHM and $\sigma_{\mathrm{noise}}$ are listed in Table 4.
Table 4: Training configuration for deCIFer and U-deCIFer.
PXRD Transformation Training Parameters Value
[…]
Q-grid ($Q_{\text{min}}$, $Q_{\text{max}}$, $Q_{\text{step}}$) (0.0, 10.0, 0.01)
FWHM $\mathcal{U}\sim(0.001,0.10)$
Mixing Factor ($\eta$) 0.5
Noise Magnitude $\mathcal{U}\sim(0.001,0.05)$
Model / Training Parameters Value
Optimizer AdamW
Learning Rate $1\times 10^{-3}$
Warmup Steps 100
Decay Steps 50,000
Minimum Learning Rate $1\times 10^{-6}$
Weight Decay 0.1
Batch Size 32
Gradient Accumulation Steps 40
Maximum Iterations 50,000
Embedding Dimension ($n_{\text{embd}}$) 512
Layers ($n_{\text{layer}}$) 8
Attention Heads ($n_{\text{head}}$) 8
Conditioning Model Layers ($n_{\text{c-layers}}$) 2
Conditioning Model Hidden Size 512
Sequence Length (block_size) 3076
Precision float16
Dropout 0.0
Section A.6 Validity Metrics
To evaluate consistency and chemical sensibility of the generated CIFs, we conduct a series of validation checks. The methodology is described below.
Formula Consistency
We check for consistency in the chemical formula printed in different locations within the CIF. Specifically, we ensure that:
  • The chemical formula in the _chemical_formula_sum tag matches the reduced chemical formula derived from the atomic sites.
  • The chemical formula in the _chemical_formula_structural tag is consistent with the composition derived from the CIF file.
Site Multiplicity Consistency
We validate that the total multiplicity of atomic sites is consistent with the stoichiometry derived from the composition. Specifically, we ensure:
  • The atom types are specified under the _atom_site_type_symbol tag.
  • The multiplicity of each atom is provided in the _atom_site_symmetry_multiplicity tag.
  • The total number of atoms derived from these tags matches the stoichiometry derived from the _chemical_formula_sum tag.
Bond Length Reasonability
To check the reasonableness of bond lengths:
  • We use a Voronoi-based nearest-neighbour algorithm implemented in the CrystalNN module of pymatgen [Ong et al., 2013] to identify bonded atoms.
  • For each bond, the expected bond length is calculated based on the atomic radii and the electronegativity difference between the bonded atoms:
    • If the electronegativity difference is greater than or equal to 1.7, the bond is treated as ionic, and the bond length is based on the cationic and anionic radii.
    • Otherwise, the bond is treated as covalent, and the bond length is based on the atomic radii.
  • A bond length reasonableness score $B$ is computed as the fraction of bonds whose lengths are within ±$\pm 30\%$ of the expected lengths.
  • A structure passes this test if $B\geq c_{\text{bond}}$, where $c_{\text{bond}}=1.0$.
Space Group Consistency
We validate the space group by:
  • Extracting the stated space group from the _symmetry_space_group_name_H-M tag.
  • Analyzing the space group symmetry using the SpacegroupAnalyzer class in pymatgen [Ong et al., 2013], which employs the spglib [Togo and Tanaka, 2018] library.
  • Comparing the stated space group with the one determined by the symmetry analysis.
Overall Validity
A CIF file is deemed valid if all the above checks are satisfied:
  • Formula consistency (FM).
  • Site multiplicity consistency (SM).
  • Bond length reasonableness $B\geq c_{\text{bond}}$, where $c_{\text{bond}}=1.0$ (BL).
  • Space group consistency (SG).
Section A.7 Match Rate
The Match Rate (MR) quantifies how many generated structures successfully match their corresponding reference structures, as determined by StructureMatcher from the pymatgen library [Ong et al., 2013]. Two structures are considered a match if their compositions, lattice parameters, atomic coordinates, and symmetry are sufficiently similar, according to the tolerances set in StructureMatcher. For the implementation of deCIFer, we follow the example set by CrystaLLM [Antunes et al., 2024], using the parameter values:
  • stol $=0.5$: site tolerance, defined as a fraction of the average free length per atom.
  • angle_tol $=10^{\circ}$: maximum angular deviation tolerance.
  • ltol $=0.3$: fractional length tolerance, meaning the lattice parameters can differ by up to 30% relative to the reference lattice.
StructureMatcher compares two structures by:
  • Optionally reducing them to primitive (Niggli) cells.
  • Verifying that the lattice parameters are within the fractional length tolerance (ltol).
  • Checking that the angles are within the angle tolerance (angle_tol).
  • Ensuring that atomic coordinates align within the site tolerance (stol), normalized by the average free length per atom.
With these parameters, each generated CIF is compared against its reference CIF*. If the two structures are deemed structurally equivalent, we count that as a successful match. MR is computed as the fraction of structures in the dataset for which a match is found:
$\mathrm{MR}\;=\;\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\bigl{(}\text{match}(\mathrm{CIF},\mathrm{CIF}^{*})\bigr{)},$ (7)
where $N$ is the total number of structures and $\mathbf{1}(\cdot)$ is an indicator function that returns $1$ if two structures match (according to StructureMatcher) and 00 otherwise.
Section A.8 Datasets Statistics
Figure 10 illustrates the NOMA dataset. Figure 11 illustrates the statistics of the curated CHILI-100K [Friis-Jensen et al., 2024] dataset.
Figure 10: Statistical overview of the NOMA [Antunes et al., 2024] dataset (2,283,346 total samples), showing the distribution of space group frequencies, the number of elements per unit cell, elemental occurrences and CIF token lengths (indicating the percentage of CIFs with larger token sequences than the context length of 3076)
Figure 11: Statistical overview of the curated CHILI-100K [Friis-Jensen et al., 2024] dataset (8201 total samples), showing the distribution of space group frequencies, the number of elements per unit cell, elemental occurrences, and CIF token lengths (indicating the percentage of CIFs with larger token sequences than the context length of 3076).
Section A.9 Model Architecture- and Training Details
Table 4 provides a concise overview of all hyperparameters and data augmentation settings used for training deCIFer (and its variant U-deCIFer). Below, we describe additional implementation details.
Data-stratification
We extract the space group number from each CIF (ranging from $1$ to $230$) and group these into bins of size ten (e.g., 111110101010, 1111111120202020, etc.). This heuristic aims to preserve the overall symmetry distribution across splits while reducing the risk of data leakage from structurally similar entries appearing in multiple subsets. While this does ensures coverage across symmetry classes and even representation of crystal systems across the splits, it does not reflect the most intuitive or principled grouping scheme. In particular, it does not account for finer-grained biases that may be embedded in crystal symmetry or composition distributions. This points to a broader issue in materials datasets: statistical artefacts, such as the "Rule of Four" or symmetry clustering [Gazzarrini et al., 2024], can introduce shortcuts that models may learn, reducing generalisation and interpretability [Palgrave, 2024]. Future work should explore alternative stratification strategies (e.g., stratified sampling based on structural descriptors) to better assess generalisation.
Hardware Setup
All experiments were conducted on GPUs with sufficient memory to accommodate a batch size of 32 tokenized sequences, each truncated or padded to a context length of 3076. We employed half-precision (float16) to reduce memory usage and improve throughput, ensuring that gradient updates remain numerically stable via built-in automatic mixed-precision.
Optimizer and Learning Rate Schedule.
We adopt AdamW with a base learning rate of $1\times 10^{-3}$, which is warmed up for 100 steps and then gradually decayed to $1\times 10^{-6}$ over 50,000 steps (Table 4). Weight decay is set to 0.1 to regularize model parameters, and we employ gradient accumulation (40 steps) to effectively increase the total number of tokens processed per update.
Transformer Architectural Notes.
The final transformer stack has 8 layers, each with 8 attention heads, and a model dimension of 512 (embedding dimension). The feed-forward blocks inside each layer use a dimension of $4\times 512$, and dropout is set to 0.0 to minimize underfitting. We continue to observe stable convergence in practice despite using no dropout.
Maximum Iterations and Convergence.
We train for 50,000 iterations, at which point the model’s cross-entropy loss stabilizes, as illustrated in Figure 12. Beyond this range, no significant improvements were observed on validation metrics.
Figure 12: Cross-entropy loss curves for U-deCIFer and deCIFer over 50,000 training iterations, showing progressive reduction in the training and validation losses.
Section A.10 PXRD Embedding Space
For completeness, we examined the learned embeddings for 50K random training-set PXRD profiles and applied principle component analysis (PCA) for visualization. As shown in Figure 13, the embeddings form distinct gradients when colored by crystal system, cell-volumes, and constituent atomic numbers $Z$, indicating that the model’s PXRD embedding captures relevant structural characteristics, such as symmetry, scale, and elemental composition. These patterns highlight the effectiveness of the conditioning mechanism in encoding meaningful structural information directly from the PXRD input.
Figure 13: 2D PCA projection of learned PXRD embeddings for 500K training-set samples from NOMA. The three subplots are colored by crystal system, $\log$(cell volume), and average atomic number $Z$, illustrating clear gradients that correspond to structural and compositional features as indicated by the arrows.
Section A.11 Baseline Comparison with and without PXRD
With regards to the baseline comparison in Table 1, deCIFer is explicitly conditioned on PXRD data, while the baseline models are conditioned only on composition or latent priors. This distinction means that the comparisons are not direct but instead reveal the relative value of PXRD conditioning. In particular, PXRD conditioning can improve structure prediction when the diffraction signal is rich in structural information, but may introduce ambiguity or conflict with the model’s learned priors in cases where the PXRD pattern is noisy or minimally informative. Perov-5 and Carbon-24 provide strong tests of PXRD conditioning due to their polyhedral complexity and carbon-based structural diversity, where diffraction features can directly inform the model. In contrast, MP-20 is drawn from the Materials Project, a dataset where composition-only models may benefit from learned priors due to the high representation of standard chemistries. MPTS-52 further challenges models with low-symmetry structures, where the PXRD signal can become ambiguous, making it difficult for a model to resolve atomic positions purely from the diffraction pattern.
Section A.12 Additional Results
Figures 16, 14, and 15 provide additional insights into deCIFer’s performance. Table 5 shows a detailsed breakdown of the validity metrics for the NOMA test set corresponding to the results in Figure 3. Table 6 shows a detailed breakdown of validity metrics for the NOMA test set and CHILI-100K test set evaluated on two in-distribution (ID) scenarios and one out-of-distribution (OOD) scenario for the PXRD input.
Section A.13 Future Work
One promising area for improvement lies in exploring more advanced decoding strategies, such as beam search, to enhance the generative model’s capabilities in downstream tasks. By maintaining multiple hypotheses during decoding, beam search could produce diverse candidate CIFs for a given PXRD profile, improving structure determination accuracy by ranking outputs based on metrics like $R_{\mathrm{wp}}$. This method could also support optimization strategies that prioritize structural validity and relevance.
Another direction could be to integrate reinforcement learning from human feedback (RLHF) to guide the model more directly toward generating accurate and chemically valid structures [Ziegler et al., 2019]. By defining a reward function tailored to properties such as low $R_{\mathrm{wp}}$ values, structural integrity, and adherence to chemical constraints, and interaction with a human expert, RLHF could further refine the model’s outputs.
A complementary direction for future improvement lies in increasing the diversity of the training data. While NOMA offers excellent scale, its distribution is skewed toward high-symmetry and well-sampled structures, which limits model generalisation to more uncommon systems. CHILI-100K, with its broader structural complexity and greater representation of low-symmetry crystals, could serve as a valuable training supplement. Mixing synthetic and experimental data, or applying curriculum learning strategies that gradually expose the model to under-represented symmetry groups, could improve generalisation and robustness, particularly for monoclinic and triclinic systems, which remain challenging as seen in Figure 4.
Table 5: Validity of generated CIFs for the NOMA test set using deCIFer and U-deCIFer. Abbreviations: Form = formula validity, SG = space group validity, BL = bond length validity, SM = site multiplicity validity. Overall validity (Val.) is calculated as the percentage of CIFs that satisfy all four validity metrics simultaneously. Match rate (MR) represents the percentage of generated CIFs that replicate the reference CIF.
Desc. Model Form (%) $\uparrow$ SG (%) $\uparrow$ BL (%) $\uparrow$ SM (%) $\uparrow$ Val. (%) $\uparrow$ MR (%) $\uparrow$
none U-deCIFer 99.82 98.87 94.30 99.47 93.49 0.00
deCIFer 99.42 98.85 93.69 99.46 92.66 5.01
comp. U-deCIFer 99.87 99.09 94.40 99.46 93.78 49.30
deCIFer 99.68 99.21 94.37 99.55 93.73 91.50
comp.+s.g. U-deCIFer 99.85 98.88 94.51 99.47 93.72 87.07
deCIFer 99.74 99.26 94.38 99.58 93.90 94.53
Table 6: Validity of generated CIFs for the CHILI-100K test set using deCIFer. Abbreviations: Form = formula validity, SG = space group validity, BL = bond length validity, SM = site multiplicity validity. Overall validity (Val.) is calculated as the percentage of CIFs that satisfy all four validity metrics simultaneously. Match rate (MR) represents the percentage of generated CIFs that replicate the reference CIF.
Dataset ($\sigma_{\mathrm{noise}}^{2}$, FWHM) FORM (%) $\uparrow$ SG (%) $\uparrow$ BL (%) $\uparrow$ SM (%) $\uparrow$ Val. (%)\;$\uparrow$ MR (%)\;$\uparrow$
NOMA ID: (0.00, 0.05) 99.68 99.21 94.37 99.55 93.73 91.50
ID: (0.05, 0.10) 99.64 99.18 94.39 99.55 93.77 89.28
OOD: (0.10, 0.20) 99.60 99.87 92.60 99.49 91.66 77.66
CHILI-100K ID: (0.00, 0.05) 95.98 97.88 42.61 94.58 41.83 37.34
ID: (0.05, 0.10) 96.17 98.22 41.50 94.47 40.95 35.97
OOD: (0.10, 0.20) 95.80 98.42 34.11 93.91 33.62 26.09
Figure 14: Average metric values by crystal systems for deCIFer on the NOMA test set under two in-distribution transformations of the input PXRD profiles and one out-of-distribution transformation. deCIFer shows better performance for well-represented systems, while rarer, low-symmetry systems lead to worse performance. The right-most plot shows crystal system distribution of the NOMA test set.
Figure 15: Average metric values by crystal systems for deCIFer on the CHILI-100K test set show better performance for well-represented systems in the training data (NOMA), while low-symmetry systems lead to worse performance. The right-most plot shows crystal system distribution of the CHILI-100K test set, highlighting that CHILI-100K contains a significantly higher proportion of lower-symmetry structures compared to synthetic datasets like NOMA.
Figure 16: Distribution of $R_{wp}$ for deCIFer on the NOMA- and CHILI-100K test set, presented as violin plots with overlain boxplots; the median is shown for each distribution. Presented are four in-distribution transformations of the input PXRD profiles and three out-of-distribution transformations.